A 5 cm diameter shaft is stepped down to 4 cm with a fillet radius of 5 mm. If the allowable shear stress of the shaft material is 55 MPa, determine the maximum torque that can be transmitted by the shaft. (Do not ignore the stress concentration factor.)



1
Expert's answer
2021-01-18T01:14:05-0500

Tmax = Ymax"\frac{(\pi)(d^3)}{16}"

= "75.35 \times \frac{(\pi)(0.4^3)}{16}\times10^6"

= 946.87 NM